In order to join a voice lesson, you pay a Php1,500.00 pesos fee, then Php500.00 for each class you go to. What is the average cost per class? Graph the function to show whether it forms a straight line or curve.
1
Expert's answer
2021-10-21T14:48:52-0400

Total cost for x number of classes = (cost per class)(x number of classes)+total fee

=500x+1500

Average cost per class "=\frac{total\space cost \space for \space x \space classes}{x}"

"y=\frac{500x+1500}{x}=500+1500/x"



Use these points to draw the graph.



So, the function is a curve.